Fermi National Laboratory
 

Operation Summary for September 22, 2010 - September 24, 2010

 

Wednesday Sept. 22

The day shift began with Operations monitoring the Tevatron (TeV) store 8106 (147.78E30) and the Recycler stash(117.0E10), with the Antiproton Source (Pbar) stacking (24.5E10), with the Linac, Booster, the and Main Injector (MI) operating normally, with the NuMI off for target work, with MiniBooNE and Meson MTest T992 and T995 experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, at 12:35 PM, Operations reported that a yard breaker for TeV sector
F0 tripped off, took out the TeV RF water system, and caused a clean abort of store 8106. Operations put the TeV through a dry squeeze. While bringing the TeV RF water system back up, Operations reported that a hose blew out in a MI quadrupole power supply (Q713) and drenched the supply. EE Support personnel began repairs and drying out of the power supply.

During the evening shift, EE Support completed their repairs of Q713 and stacking resumed by 5:54 PM. Operations disabled beam to MTest at 6:30 PM. TeV personnel performed parsing studies until 8:50 PM. Operators replaced a fuse for a TeV power supply (T:VE47) regulator. Operations put the TeV into HEP shot setup at 10:15 PM, however the shot setup was delayed due to bad beam positions at F1 when extracting beam into the A1 transfer line. Operators investigated and thought the cause was due to the Q713 power supply.

Thursday Sept. 23

The midnight shift began with operators investigating Q713, with Operations monitoring the stash (326.8E10), with Pbar stacking (41.2E10), and with MiniBooNE taking beam.

During the midnight shift, operators called in EE Support personnel to help investigate Q713. They discovered a problem with the PLC communication protocol. Once this was reloaded, the power supply functioned properly. Operations put the TeV into shot setup at 6:35 AM and established store 8112 at 7:43 AM with an average initial luminosity (AIL) of 299E30.

During the day shift, Operations established beam to T992 and T995 at 8:15 AM. Operations halted the MI ramp so that EE Support personnel could look at the MI-10 quadrupole power supply.

During the evening shift, Operations disabled beam to T992 and T995 at 6:05 PM. TeV personnel began end-of-store studies at 7 PM. Operations terminated store 8112 at 9 PM, put the TeV into shot setup, and established store 8113 at 10:30 PM with an AIL of 298E30.

Friday Sept. 25

The midnight shift began with Operations monitoring store 8113 (201.13E30) and the stash (86.2E10), with Pbar stacking (17.4E10), and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8113 at 7:35 AM. Operations turned off the TeV for scheduled LCW system repairs at TeV sectors C0 and DA, and for an A-E access to address vacuum issues. CDF and D0 personnel also accessed their collision halls.

The Plans for Friday and the Weekend

The plans for today are to complete the TeV LCW system repairs and then resume stacking, stashing, and storing for the rest of today and the weekend.

 

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/27/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory